OpenStudy (anonymous):

PLEASE HELP!!!!!!! Choose the ordered pair that is a solution to the system of equations. y - x = 5 2y+x=-2 A. (-4,-1) B. (-1,-4) C. (-4, 1) D. (4, -1)

OpenStudy (anonymous):

C

OpenStudy (anonymous):

1--4=5 and 2(1)+-4=-2

OpenStudy (anonymous):

y-x = 5 2y + x = -2 -------------- multiply top line by -2 to get: -2y +2x = -10 2y + x = -2 -------------- 0 +3x = -12 x = -4 substituting x in the first equation (i.e. y- x =5) we get: y - (-4) = 5 y + 4 = 5 y = 5 -4 y = 1 therefore solution is: (-4, 1) which is C